30,000 people attended a ballgame at a stadium that offers two kinds of seats: general admission and reserved. The day's receipts were $201,000. How many people paid $13.00 for reserved seats, and how many paid $4.00 for general admission?
Answer by checkley77(12844)

X+Y=30,000
X=30,000-Y
13X+4Y=201,000
13(30,000-Y)+4Y=201,000
390,000-13Y+4Y=201,000
-9Y=201,000-390,000
-9Y=189,000
Y=189,000/9
Y=21,000 $4.00 TICKETS WERE SOLD.
30,000-21,000=9,000 $13.00 TICKETS WERE SOLD.
PROOF:
13*9,000+4*21,000=201,000
117,000+84,000=201,000
201,000=201,000